OpenStudy (anonymous):

in triangle ABC, A=32 degrees, a=8, and b=7. Find C. A. 110.7 degrees B. 20.7 degrees C. 30.4 degrees D. 120.4 degrees

OpenStudy (anonymous):

what have you done so far?

OpenStudy (anonymous):

what you mean?

OpenStudy (anonymous):

to do this problem.... did you use the Law of Sines?

OpenStudy (anonymous):

im not sure how to do this problem.. ha i havent done anything. thats the question

OpenStudy (anonymous):

 this is a problem using the Law of Sines.... \(\large \frac{sinA}{a}=\frac{sinB}{b}=\frac{sinC}{c} \) can you apply it?